# Adapting Multi-modal Large Language Model to Concept Drift in the Long-tailed Open World

### v1

@@ Abstract
Real-world data often exhibit extreme imbalances and out-of-distribution (OOD) instances, which significantly biases the model training. While it has been extensively studied in vision and language domains separately, the impact of long-tailed open worlds on multi-modal large language models (MLLMs) has been largely overlooked. In this paper, we first demonstrate the susceptibility and vulnerability of vision-language models to significant biases caused by tail drift and out-of-distribution (OOD) drift during both the pre-training and fine-tuning stages. To eliminate the bias from different sources, we integrate the tailed drift adaptation and OOD drift detection into a unified framework by extending the concept drift theory to multi-modal. Specifically, a T-distribution-based drift adapter is proposed to effectively mitigate the bias induced by the long-tailed problem, which also facilitates the model in distinguishing OOD data through explicit distribution modelling. Extensive experiments show significant improvements in our model’s ability to adapt to tailed drift and OOD drift. Moreover, it enhances the efficiency and accuracy of image-text alignment in vision language model pre-training, particularly in the long-tail open world scenario. Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open world scenario, to validate our findings. To foster the development of the multi-modal community, we have made both OpenMMlo datasets and our code publicly available at: https://github.com/Anonymous0Knight/ConceptDriftMLLMs.
@@ 1 Introduction
The explosive growth of data has precipitated a crisis for multi-modal large language models (MLLMs), which struggle to adapt to the long-tailed open world, encompassing the tailed drift and the out-of-distribution (OOD) drift. The tailed drift is brought by extremely imbalanced data, where the distribution of tailed categories shifts due to their inherent sparsity and noise, as well as the dominant head categories’ squeezing effect in the feature space. Meanwhile, OOD drift often occurs alongside the tailed drift and disrupts the underlying distribution of the model, as we continually encounter new concepts while navigating in an open world. Although the critical challenge of the long-tailed open world has been increasingly drawing attention in the realms of visual models [1] and language models [2], its impact on MLLMs, particularly vision-language (VL) models, has received scant attention. Therefore, we provide a comprehensive analysis of the impact of tail drift and OOD drift on VL models during both the pre-training and fine-tuning phases.
@@
Therefore, we propose a concept drift-aware multi-modal large language model, mitigating the tail drift and OOD drift encountered in the long-tailed open world. Firstly, we introduce the concept drift theory to the multi-modal domain, which provides a more holistic perspective to explain tailed drift and OOD drift. Then, the T-distributed adapter is proposed to be embedded in the hyperspherical feature space. It aligns image-text features for contrastive learning in pre-training. The desirable light-tailed property of the proposed T-distributed spherical metric (tSP) prevents the compression of tailed categories and mitigates the crowding of feature space caused by tailed drift. Besides, in fine-tuning, the adapter projects features to decision boundary and detects OOD samples at feature-level based on the underlying distribution. The proposed tSP distribution explicitly models the feature space with concrete feature centers, optimizing large inter-class margins and yielding more desirable hyperspherical embeddings. And a simple non-parametric KNN is adopted to distinguish the OOD sample based on the tSP distribution. Finally, we construct a group of multi-modal long-tailed open datasets to support our claims.
@@
    We are the pioneers in revealing the unexplored impacts of long-tailed open world to multi-modal large language models, especially in the image-text alignment in the pre-training and feature space allocation in the fine-tuning. This allows future research to more comprehensively study the impact of defect data on MLLMs.
@@ 2.1 Multi-modal Concept Drift Theory
Assume that there are $N$ data streams corresponding to $N$ modalities, given a set of examples denoted as $S_{0,t}=\{S_{0},...,S_{i},...,S_{t}\}$, where $S_{i}=(s_{1},...,s_{j},...,s_{N})$ and $s_{j}=(X_{ij},y_{i})$ is one data instance from a single $j$-th data stream, $X_{ij}$ is the feature vector, $y_{i}$ is the label and $t$ is the timestamp of the data stream. $S_{0,t}$ follows a certain distribution $F_{0,t}(S_{i})$, the multi-modal concept drift occurs at timestamp $t+1$, if $F_{0,t}(S_{i})\neq F_{t+1,\infty}(S_{i})$, denoted as:
@@
2.2 T-distributed Adapter for Long-Tailed Open World
Within the concept drift framework, data distribution-based drift detection and adaptation have emerged as the dominant approaches [5, 6]. To effectively mitigate the concept drift in the long-tail open world, we integrate directional statistics into distribution-based drift detection and adaptation, and propose a T-distributed adapter to alleviate it in the vision language model.
2.2.1 Directional statistics
Directional statistics primarily focus on the distribution of eigenvector angles, while neglecting the impact of eigenvector module lengths. Given the unit feature vector $X_{ij}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:||x||_{2}=1\}$ denotes the $(d-1)$-dimensional hyperspherical set. A key idea in directional distribution is the tangent-normal decomposition. Any unit vector $x$ can be decomposed as:
$x=t\mu+(1-t^{2})^{\frac{1}{2}}v,t\in[-1,1]$ (2)
with $v\in\mathbb{S}^{d-2}$ a tangent to $\mathbb{S}^{d-1}$ at $\mu$ [7, 8], where $v$ and $t$ are independent and $v$ is uniform on $\mathbb{S}^{d-2}$. Thus, the intersection of $\mathbb{S}^{d-1}$ with the hyperplane through $t\mu$ and normal to $\mu$ is a $(d-2)$-dimensional sphere of radius $\sqrt{1-t^{2}}$, that $t$ has density as following:
$p_{T}(t;d)\propto(1-t^{2})^{\frac{d-3}{2}},t\in[-1,1]$ (3)
Therefore, through the marginal density $p_{T}$ and $p_{v}$, we can estimate the density of the entire spherical distribution. One prominent instance is the von Mises-Fisher distribution (vMF) [9], which can be interpreted as a probability distribution over the cosine similarity between a unit vector $x$ and a fixed mean direction $\mu$, following the density:
$p_{X}(x;\mu,\kappa)\propto\exp{(\kappa\mu^{T}x)}$ (4)
where $\kappa\geqslant 0$ denotes the concentration and $\exp$ represents the exponential function. Therefore, combined with the Eq. 2 and Eq. 3, the density of vMF is:
$p(x)=C_{X}(\kappa,d)^{-1}\exp{(\kappa\mu^{T}x)},\quad x\sim\text{vMF}(\mu,\kappa)$ (5)
$C_{X}(\kappa,d)=\frac{(2\pi)^{d/2}I_{d/2-1}(\kappa)}{\kappa^{d/2-1}}$
where $I_{m}$ denotes the modified Bessel function of the first kind at order $m$.
2.2.2 T-distributed Adapter
Although the spherical nature of the vMF mitigates the bias induced by module length in long-tail open worlds, its exponential property, as evident in Eq. 4, leads to a winner-take-all behaviour, where only a marginally higher logit than its competitors will dominate the eigenspace, receiving an exponentially disproportionate share of the eigenspace. As a consequence, the head categories, which possess abundant samples, exert a dominant influence over the entire model and squeeze the eigenspace of tailed categories, causing the crowding problem.
@@ 2.2.2 T-distributed Adapter
Therefore, inspired by the T-SNE [10], we design a T-distribution based spherical metric (tSP), which follows the density:
$p_{X}(x)\propto\frac{2}{\kappa(1-\mu^{T}x)}$ (6)
where the unit feature vector $x\in\mathbb{S}^{d-1}$ and the center $\mu\in\mathbb{S}^{d-1}$. Through Eq. 3, we can get the marginal normalizer:
$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (7)
@@
where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. 2, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
@@
Thus, the probability density function of the proposed tSP distribution is as follows:
$p(x)=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x)},\quad x\sim\text{tSP}(\mu)$ (9)
The detailed derivation process is provided in Appendix A.2.
@@
In terms of adapting to tailed concept drift, the tSP metric with a large concentration exhibits a light-tailed property, wherein the probability density function exhibits a faster rate of decay as the values increase, relative to the vMF metric, as illustrated in Figure 2. The high kurtosis of tSP is characterized that it yields high confidence only when the feature vector is sufficiently close to the center of the category, thereby minimizing the influence of head category samples on the tail category centers. Formally, $L_{\text{tsp}}(\mu,x)=\frac{2}{\kappa(1-\mu^{T}x)+\epsilon}$ denotes the T-distributed spherical metric where $\epsilon$ is a non-zero value setting to $1$ to avoid the denominator of 0, and $L_{\text{vmf}}(\mu,x)=\exp(\kappa\mu^{T}{x})$ represents the vMF metric. Given an unit feature vector $x_{\text{head}}$ from the head categories, the gradient of the metric $L$ over the tailed category center $\mu_{\text{tail}}$ is $\frac{\partial L(\mu_{\text{tail}},x_{\text{head}})}{\partial\mu_{\text{tail}}}$. Due to $\mu^{T}_{\text{tail}}x_{\text{head}}\in[-1,1]$, when $\kappa\geqslant 1$, it is readily obtain that $\frac{\partial L_{\text{tsp}}}{\partial\mu_{\text{tail}}}<\frac{\partial L_{\text{vmf}}}{\partial\mu_{\text{tail}}}$. Consequently, the light-tailed tSP distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by tailed concept drift.
Likewise, the tSP metric is directly applied to detect the OOD concept drift. A sample with a unit feature vector $x$ is deemed out-of-distribution if it lies at a relatively large distance from the in-distribution (ID) data in the eigenspace. Following [11], a simple non-parametric KNN is adopted to partition the data into two sets (ID vs. OOD), which does not impose any distributional assumption on the feature space. Here the distance is the tSP metric with respect to the k-th nearest neighbor.
Figure 3: The workflow of our methodology, consisting of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks. By framing multi-modal heterogeneous data as concept drift data streams, we unify both tail drift and OOD drift. In the pre-training, the tSP metric aligns visual and textual feature space by image-text contrastive learning, with a large inter-class margin. Coupled with the language model loss, they drive the training of all the module. In the downstream task, the image encoder and the text decoder are frozen out of training, and two linear layers as the projector embed features into the hyperspherical eigenspace. The T-distributed adapter explicitly incorporates light-tail property into the feature space modelling, allowing it to effectively adapt tail drift, and performs OOD drift detection based on the distribution.
2.3 T-distributed Vision Language Model for the Long-Tailed Open World
As illustrated in Figure 3, our proposed vision language model contains two stages, the pre-training and the fine-tuning on the downstream task. Specifically, the classification task is chosen to visualize the impact of the bias caused by the long-tailed open world on the model. The multi-modal concept drift theory offers a unified framework for integrating tail drift adaptation and OOD drift detection, where heterogeneous image and text inputs are treated as distinct data streams.
@@ 2.3 T-distributed Vision Language Model for the Long-Tailed Open World
In the downstream classification task, we leverage the T-distributed adapter to mitigate the long-tailed distribution of categories, thereby improving classification performance. Furthermore, by explicitly modelling the feature space, the T-distributed adapter enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples. Besides, following the Blip [15], the image encoder and the text decoder are frozen out of training during fine-tuning. A head of the classifier with two linear layers embeds features into the spherical eigenspace is trained.
2.4 Building Multi-modal Datasets Group OpenMMlo for the Long-Tailed Open World
@@
We extend the open-source datasets, namely ImageNet-LT [16], iNatualist2018 [17] and Places-LT [16]. ImageNet-LT has 1,000 classes and contains 115.8k samples, with a maximum of 1,280 samples and a minimum of 5 samples for a category. Besides, it consists of 18k images for OOD detection. Places-LT has 184.5K samples from 365 classes, with class samples ranging from 4,980 to 5. The iNaturalist 2018 is a large-scale species dataset collected in the natural world with 437.5K samples for 8,142 classes. We use the InstructBLIP [18] to generate the related caption of the image, with the prompt of "What does this picture describe? Please describe in detail its size, location, color, and its relationship to the surroundings.". For more details about OpenMMlo, please refer to Appendix A.4.
@@ 3 Experiments
In this section, we first present the performance in downstream long-tailed classification and OOD detection tasks, which is induced by tail drift and OOD drift on MLLMs. Then, we evaluate the interior feature space of the VL model and further demonstrate our method alleviates the crowding and bias problems caused by the tail drift and OOD drift. The constructed long-tailed multi-modal datasets OpenMMlo is utilized for training and validating. In terms of the OOD drift detection, we follow the setting of CIDER [19]. The model is trained on CIFAR100-LT [20] with an imbalance ratio of 100, and validated on external OOD datasets including SVHN [21], Places365 [22], LSUN [23], iSUN [24] and Texture [25]. More detailed experimental implementations are given in Appendix A.3.

### v2

# Adapting Multi-modal Large Language Model to Concept Drift From Pre-training Onwards

@@ Abstract
Multi-modal Large Language Models (MLLMs) frequently face challenges from concept drift when dealing with real-world streaming data, wherein distributions change unpredictably. This mainly includes gradual drift due to long-tailed data and sudden drift from Out-Of-Distribution (OOD) data, both of which have increasingly drawn the attention of the research community. While these issues have been extensively studied in the individual domain of vision or language, their impacts on MLLMs in concept drift settings remain largely underexplored. In this paper, we reveal the susceptibility and vulnerability of Vision-Language (VL) models to significant biases arising from gradual drift and sudden drift, particularly in the pre-training. To effectively address these challenges, we propose a unified framework that extends concept drift theory to the multi-modal domain, enhancing the adaptability of the VL model to the distribution unpredictable changes. Additionally, a T-distribution based drift adapter is proposed to effectively mitigate the bias induced by the gradual drift, which also facilitates the model in distinguishing sudden distribution changes through explicit distribution modeling. Extensive experiments demonstrate our method enhances the efficiency and accuracy of image-text alignment in the pre-training of VL models, particularly in the concept drift scenario. Moreover, various downstream tasks exhibit significant improvements in our model’s ability to adapt to long-tailed open world. Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open world settings, to validate our findings. To foster the development of the multi-modal community, we have made both OpenMMlo datasets and our code publicly available at: https://github.com/Anonymous0Knight/ConceptDriftMLLMs.
@@ 1 Introduction
The rapid expansion of data availability has created significant challenges for multi-modal large language models (MLLMs), particularly in addressing concept drift, which predominantly manifests as gradual drift and sudden drift [1]. Among them, tailed drift represents a classic illustration of gradual drift, emerging due to severe data imbalance, where the distributions of long-tail categories evolves because of their intrinsic sparsity and noise. Concurrently, sudden drift is mainly represented by OOD drift, as the model encounters new, previously unseen concepts, resulting in distributional shifts that disrupt its ability to generalize in an open-world context. While the issues of long-tailed recognition and concept drift in open-world settings have been extensively studied in visual models [2] and language models [3], their impact on MLLMs, particularly vision-language (VL) models, remains largely unexplored. In this work, we aim to bridge this gap by providing a systematic analysis of how tailed drift and OOD drift affect VL models during both pre-training and fine-tuning phases. Our findings highlight critical vulnerabilities of current VL models in adapting to these challenges, underscoring the need for novel strategies to enhance their robustness in dynamic, open-world environments.
@@
Therefore, we propose a concept drift-aware multi-modal large language model, mitigating the tail drift and OOD drift encountered in the long-tailed open world. Firstly, we introduce the concept drift theory to the multi-modal domain, which provides a more holistic perspective to explain tailed drift and OOD drift. Then, the T-distributed adapter is proposed to be embedded in the hyperspherical feature space. It aligns image-text features for contrastive learning in pre-training. The desirable light-tailed property of the proposed T-distributed spherical metric (Thp) prevents the compression of tailed categories and mitigates the crowding of feature space caused by tailed drift. Besides, in fine-tuning, the adapter projects features to decision boundary and detects OOD samples at feature-level based on the underlying distribution. The proposed T-hp distribution explicitly models the feature space with concrete feature centers, optimizing large inter-class margins and yielding more desirable hyperspherical embeddings. And a simple non-parametric KNN is adopted to distinguish the OOD sample based on the T-hp distribution. Finally, we construct a group of multi-modal long-tailed open datasets to support our claims.
@@
    We are the pioneers in revealing the unexplored impacts of concept drift to multi-modal large language models, especially in the image-text alignment in the pre-training and feature space allocation in the fine-tuning. This allows future research to more comprehensively study the impact of defect data on MLLMs.
@@ 2.1 Multi-modal Concept Drift Theory
Assume that there are $N$ data streams corresponding to $N$ modalities, given a set of examples denoted as $S_{0,t}=\{S_{0},...,S_{i},...,S_{t}\}$, where $S_{i}=(s_{1},...,s_{j},...,s_{N})$ and $s_{j}=(X_{ij},y_{i})$ is one data instance from a single $j$-th data stream, $X_{ij}$ is the feature vector, $y_{i}$ is the label and $t$ is the timestamp of the data stream. $S_{0,t}$ follows a certain distribution $F_{0,t}(S_{i})$, the multi-modal concept drift occurs at timestamp $t+1$, if $P_{0,t}(S_{i})\neq P_{t+1,\infty}(S_{i})$, denoted as:
@@
Figure 2: The workflow of our methodology, consisting of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks. Within the data streaming, a drift adaptation window slides to detect changes in data distribution and subsequently update the model, in both pre-training and fine-tuning. In the pre-training, the T-distributed adapter aligns visual and textual feature space by image-text contrastive learning, with a large inter-class margin. Coupled with the language model loss, they drive the training of all modules. In the downstream task, the image encoder and the text decoder are frozen out of training, with a linear projector fusing image-text features. Additionally, a mixture of experts module is leveraged with the T-distributed adapter as the router, allowing it to effectively adapt tail drift and perform OOD drift detection based on the distribution.
2.2 T-distributed Adapter for Concept Drift
To adapt the vision language model to concept drift, it is essential to adapt the model to align with the evolving data distribution, which can be formally defined as
$\min_{f^{(t)},f^{(t+1)},...,f^{(t+\tau)}}\sum_{i=t}^{t+\tau}L(f^{(i)}(x^{(i)}),y^{(i)})$ (2)
where $f^{(t)}$ denotes the vision language model trained by the data stream $S_{t-k,t-1}$ from the drift adaption window with the size of $k$. And the model is driven by the target metric $L$ continuously to adapt the drift in a given time period $[t,t+\tau]$. Thus, one prevalent method for detecting and adapting to concept drift involves designing metrics based on data distributions that can effectively counteract the impacts of sudden and gradual changes within the time window [5, 6]. Building upon this thinking, we integrate directional statistics into distribution-based drift detection and adaptation, proposing a T-distributed adapter to alleviate it in the vision language model. Firstly, we provide an overview of directional statistics in the Appendix LABEL:appendix:ds. Then, we will introduce the T-distributed adapter.
@@ 2.2 T-distributed Adapter for Concept Drift
Inspired by the T-SNE [7], we design a T-distribution based metric in hypersphere (T-hp), which follows the density:
$p_{X}(x^{(i)})\propto\frac{2}{\kappa(1-\mu^{T}x^{(i)})}$ (3)
where $x^{(i)}\in\mathbb{S}^{d-1}$ denotes the unit feature vector, $\mu\in\mathbb{S}^{d-1}$ represents the center of category and $\kappa\geq 0$ symbolizes the concentration of the distribution, with higher values indicating a greater concentration around the center $\mu$. Accordingly, we can get the marginal normalizer:
$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x^{(i)}}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (4)
@@
where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. LABEL:eq:2, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
@@
Thus, the probability density function of the proposed Thp distribution is as follows:
$p(x^{(i)})=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x^{(i)})},\quad x^{(i)}\sim\text{Thp}(\mu)$ (6)
The detailed derivation process is provided in Appendix LABEL:appendix:Tsp.
@@
In terms of adapting to tailed concept drift, the Thp metric with a large concentration exhibits a light-tailed property, wherein the probability density function exhibits a faster rate of decay as the values increase, relative to the vMF metric, as illustrated in Figure 3. The high kurtosis of Thp is characterized that it yields high confidence only when the feature vector is sufficiently close to the center of the category, thereby minimizing the influence of head category samples on the tail category centers. Formally, $L_{\text{thp}}(\mu,x^{(i)})=\frac{2}{\kappa(1-\mu^{T}x^{(i)})+\epsilon}$ denotes the T-distributed metric on hypersphere, where $\epsilon$ is a non-zero value setting to $1$ to avoid the denominator of 0, and $L_{\text{vmf}}(\mu,x)=\exp(\kappa\mu^{T}{x})$ represents the vMF metric. Given an unit feature vector $x_{\text{head}}$ from the head categories, the gradient of the metric $L$ over the tailed category center $\mu_{\text{tail}}$ is $\frac{\partial L(\mu_{\text{tail}},x_{\text{head}})}{\partial\mu_{\text{tail}}}$. Due to $\mu^{T}_{\text{tail}}x_{\text{head}}\in[-1,1]$, when $\kappa\geqslant 1$, it is readily obtain that $\frac{\partial L_{\text{thp}}}{\partial\mu_{\text{tail}}}<\frac{\partial L_{\text{vmf}}}{\partial\mu_{\text{tail}}}$. Consequently, the light-tailed Thp distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by tailed concept drift.
Likewise, the Thp metric is directly applied to detect the OOD concept drift. A sample with a unit feature vector $x$ is deemed out-of-distribution if it lies at a relatively large distance from the in-distribution (ID) data in the eigenspace. Following [8], a simple non-parametric KNN is adopted to partition the data into two sets (ID vs. OOD), which does not impose any distributional assumption on the feature space. Here the distance is the Thp metric with respect to the k-th nearest neighbor.
2.3 T-distributed Vision Language Model for the Concept Drift
As illustrated in Figure 2, our proposed vision language model contains two stages, the pre-training and the fine-tuning on the downstream task. Specifically, the classification task is chosen to visualize the impact of the bias caused by the long-tailed open world on the model. The multi-modal concept drift theory offers a unified framework for integrating gradual drift adaptation and sudden drift detection, where heterogeneous image and text inputs are treated as distinct data streams.
@@ 2.3 T-distributed Vision Language Model for the Concept Drift
In the downstream classification task, we leverage the T-distributed adapter as the router to distribute features to various FFNs as experts. Furthermore, by explicitly modelling the feature space, the T-router enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples. Besides, following the Blip [12], the image encoder and the text decoder are frozen out of training during fine-tuning. A head of the classifier with two linear layers embeds features into the spherical eigenspace is trained.
2.4 Building Multi-modal Dataset OpenMMlo for the Long-Tailed Open World
@@
We extend the open-source datasets, namely ImageNet-LT [13], iNatualist2018 [14] and Places-LT [13]. ImageNet-LT has 1,000 classes and contains 115.8k samples, with a maximum of 1,280 samples and a minimum of 5 samples for a category. Besides, it consists of 18k images for OOD detection. Places-LT has 184.5K samples from 365 classes, with class samples ranging from 4,980 to 5. The iNaturalist 2018 is a large-scale species dataset collected in the natural world with 437.5K samples for 8,142 classes. We use the InstructBLIP [15] to generate the related caption of the image, with the prompt of "What does this picture describe? Please describe in detail its size, location, color, and its relationship to the surroundings.". For more details about OpenMMlo, please refer to Appendix LABEL:appendix:details.
@@ 3 Experiments
In this section, we first present the performance in downstream long-tailed classification and OOD detection tasks, which is induced by tail drift and OOD drift on MLLMs. Then, we evaluate the interior feature space of the VL model and further demonstrate our method alleviates the crowding and bias problems caused by the tail drift and OOD drift. The constructed long-tailed multi-modal datasets OpenMMlo is utilized for training and validating. In terms of the OOD drift detection, we follow the setting of CIDER [16]. The model is trained on CIFAR100-LT [17] with an imbalance ratio of 100, and validated on external OOD datasets including SVHN [18], Places365 [19], LSUN [20], iSUN [21] and Texture [22]. More detailed experimental implementations are given in Appendix LABEL:appendix:expdetails.